OpenStudy (anonymous):

Sand is falling on a pile , having shape pf a cone, at the rate 1.5 m3/min. assime diameter of base is 3 times the altitiude, at what rate is the altitude increasing when the altitude is 2m?

OpenStudy (anonymous):

@ganeshie8 @ParthKohli @Loser66

OpenStudy (anonymous):

whats is "r" equal too ?

OpenStudy (anonymous):

the question says 3d=h

OpenStudy (anonymous):

radius is half so 2/3r=h?

OpenStudy (anonymous):

other way around. d = 3h

OpenStudy (anonymous):

so r = 3h/2

OpenStudy (anonymous):

and V = πr²h/3, so make the substitution for r then take the derivative

OpenStudy (anonymous):

okey i did eberyhing i am getting 1/9pi and u ?

OpenStudy (anonymous):

i did somthing wrong

OpenStudy (anonymous):

yeah I got something different? What was your derivative?

OpenStudy (anonymous):

did u get 1/6pi?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

If you mean \[\Large \frac{1}{6\pi}\] then you are correct

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

oh, that's what I thought you meant

jimthompson5910 (jim_thompson5910):

what did you get for dV/dt ?

OpenStudy (anonymous):

my der was 9/12pih3 dh/dt?

OpenStudy (anonymous):

after putiing in 3/2h

jimthompson5910 (jim_thompson5910):

the h^3 part is incorrect

OpenStudy (anonymous):

how

jimthompson5910 (jim_thompson5910):

you will have h^3 in the V volume formula but after you derive, it turns into 3h^2

OpenStudy (anonymous):

the formula for cone is x=1/3r2h

OpenStudy (anonymous):

when 3/2h goes in it become h2 then h3 by timising the outside h

OpenStudy (anonymous):